OpenStudy (anonymous):

Help : Determine if y=2x+1 is a function.

OpenStudy (amistre64):

how do you determine it?

OpenStudy (anonymous):

yea it is .....

OpenStudy (akshay_budhkar):

Yes it is very much a function? What conditions are required for determining?

OpenStudy (angela210793):

if every x has a different y then yes it is

OpenStudy (amistre64):

angela; that is not quite right, might be lost in translation tho

OpenStudy (anonymous):

Btw Thank you all for jumping in to help me I do not see this happen to every ones problem :P

OpenStudy (amistre64):

if every x value produces a single y value

OpenStudy (akshay_budhkar):

amistre? why? if a a x produces two y values then???

OpenStudy (anonymous):

Crap is it ok to give a good answer to every one who helped? :S they are all good :P

OpenStudy (angela210793):

hmmmm...tht's wht i meant Amistre :)

OpenStudy (anonymous):

the actual function is \[\{(x,y)|y = 2x+1\}\]

OpenStudy (amistre64):

a function is defined as a special relation in which the dependant variable can be mapped onto only one range element. So that when we are trying to determine the results of a given model we can predict an outcome

OpenStudy (anonymous):

It is a copy of the problem i can send pic if you like?

OpenStudy (anonymous):

if you want a function you really have to write it as one. so we say \[y=2x+1\] DETERMINES a function, not IS a function

OpenStudy (amistre64):

y^2 = x+2; can be graphed, and if we use a conventional system of a vertical y axis and a horizontal x axis; each x value produces 2 results

OpenStudy (amistre64):

determines a function is good :)

OpenStudy (anonymous):

A function is a subset of a Cartesian (e.g. \(\mathbb{R}\times\mathbb{R}\) )product such that, if (x,a) and (x,b) are in the subset then a=b

OpenStudy (akshay_budhkar):

so a relation in which one x produces two y's is not a function? then what is it called? for eg x^2=y?

OpenStudy (amistre64):

x^2 =y can be considered a function; if we use conventional graphing means

OpenStudy (akshay_budhkar):

y^2=x i mean

OpenStudy (anonymous):

what do you mean by conventional ?

OpenStudy (amistre64):

x=y^r is only a funtion if r=1

OpenStudy (anonymous):

I think i got lost lol

OpenStudy (amistre64):

the standard graphing system denotes y as a vertical axis and x as a horizontal axis

OpenStudy (amistre64):

if those conventions are not kept than all bets are off :)

OpenStudy (anonymous):

got it : )

OpenStudy (anonymous):

that helped :D

OpenStudy (akshay_budhkar):

but we can plot y^2=x on a graph using conventions too no?

OpenStudy (amistre64):

plotting is a result of being a "relation". A function is a "special type of relation"

OpenStudy (akshay_budhkar):

So are these relations classified further? and what are the classes then?

OpenStudy (amistre64):

and a one to one relation is a special type of "function" lol

OpenStudy (akshay_budhkar):

one is to many is a function? but many is to one is not ight?

OpenStudy (anonymous):

wow thank you so much for helping me under stand it :D i love how there are like 15 people looking at my problem :D

OpenStudy (amistre64):

the whole gambit of classes or relations, i dont think I know of; just the basic few

OpenStudy (anonymous):

It should be noted that a graph is a function. In fact you are just setting up a injective map between the function (set of ordered pairs) and \(\mathbb{R}^2\) the vector space.

OpenStudy (amistre64):

one to one is a function that has an inverse many to one can be a function one to many is a relation

OpenStudy (anonymous):

ok i have a question how is function determined in 3d space

OpenStudy (akshay_budhkar):

ya exactly! How can we talk of function in 3 dimensions>?

OpenStudy (amistre64):

easy; when i swing a hammer in 3d space; the function is to drive a nail :)

OpenStudy (akshay_budhkar):

right! but what are the conditions that will apply? like for 2 dimensions?

OpenStudy (anonymous):

but how do you say that it is function

OpenStudy (akshay_budhkar):

we can also take angular displacement or angular velocity as a function in 3 dimension.

OpenStudy (amistre64):

as long as what your mapping and what your mapping onto conforms to the standards; its good

OpenStudy (anonymous):

\[f:\mathbb{R}\to\mathbb{R}^3\]\[f:\mathbb{R^3}\to\mathbb{R}\]\[f:\mathbb{R^3}\to\mathbb{R^3}\] Do you see now talking about functions in 3-space is ambiguous.

OpenStudy (akshay_budhkar):

In fact you are just setting up a injective map between the function (set of ordered pairs) and R2 the vector space. I couldnt get this alchemista.

OpenStudy (akshay_budhkar):

What does that mean? r^2?? r means Real numbers?

OpenStudy (anonymous):

When you raise a field to a power you are taking the Cartesian product of the field

OpenStudy (anonymous):

\[\mathbb{R}^2 = \mathbb{R}\times \mathbb{R}\]

OpenStudy (akshay_budhkar):

R^3 means you are taking the cartesian product twice?

OpenStudy (anonymous):

\[\mathbb{R}^3 = \mathbb{R} \times \mathbb{R} \times \mathbb{R}\]

OpenStudy (anonymous):

that would be an (x, y, z), where x, y, and z are real numbers.

OpenStudy (akshay_budhkar):

f:R→R3 f:R3→R f:R3→R3 Do you see now talking about functions in 3-space is ambiguous. why is it ambigious then?

OpenStudy (anonymous):

I said it was ambiguous because when you talk about functions in 3space it could mean more than one thing.

OpenStudy (anonymous):

From which space to which space?

OpenStudy (anonymous):

This is good stuff right here xD

OpenStudy (akshay_budhkar):

Can you give an example alchemista?

OpenStudy (anonymous):

I just showed you above, those three examples you pasted again.

OpenStudy (akshay_budhkar):

no no an example can you explain by taking a function?

OpenStudy (anonymous):

Oh err someone else think of a real world example

OpenStudy (anonymous):

What do you want an example of? a function from R^3 to R^3? or R^3 to R? or whatever combination? lol

OpenStudy (anonymous):

sphere is it a function in 3d space

OpenStudy (akshay_budhkar):

A combination which shows that three dimensional function can be ambiguous?

OpenStudy (anonymous):

a sphere is a set of points in R^3 that satisfy some condition.

OpenStudy (anonymous):

No, I said that when you said a 3d function, it doesn't tell us what you mean. Not that a function is ambiguous.

OpenStudy (akshay_budhkar):

yes a sphere can be a good example

OpenStudy (anonymous):

i think i should read and then talk with you guys sorry for the trouble

OpenStudy (anonymous):

I think what alchemista is saying is that if you just say "a 3d function", its ambiguous because we dont know what the inputs and outputs are.

OpenStudy (akshay_budhkar):

See final query here Like a relation in a 2 dimensional convention be not a function can similarly a case in 3 dimensional relation occure? Yes or no plz give 1 eg if possible.. yes ishaan i feel the Same now

OpenStudy (anonymous):

You have to be more specific, like, "a function from R to R^3" that tells us it takes in one dimensional objects and spits out 3 dimensional objects

OpenStudy (anonymous):

Yes joemath that's what I mean

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Three-dimensional_graph look at his so unusual ..i never saw something like that

OpenStudy (akshay_budhkar):

ok say a function from r^3 to R^3

OpenStudy (akshay_budhkar):

Give me one eg of such a function

OpenStudy (anonymous):

T(x,y,z) = (x+y, y+z, x+z)

OpenStudy (anonymous):

in theory you have to do that anyway. to specify a function you are supposed to define the domain and range no matter what dimension you are in

OpenStudy (anonymous):

thats true satellite.

OpenStudy (akshay_budhkar):

why is it r^3 to r^3 but? I am not getting that!

OpenStudy (anonymous):

which is why math teachers should stop with problems that say "find the domain of..."

OpenStudy (anonymous):

Lets just start with linear transformations

OpenStudy (anonymous):

Something from R^3->R^3 could take some point in 3space and rotate it for instance.

OpenStudy (anonymous):

its R^3 to R^3 because your input is 3 tuple (did i say that right?)

OpenStudy (anonymous):

You can't really graph that.

OpenStudy (anonymous):

yeah =/

OpenStudy (anonymous):

sorry i didn't knew that

OpenStudy (akshay_budhkar):

yes alchemista.. i took a point in 3-d n rotated it then??

OpenStudy (akshay_budhkar):

Oh i think i am taking it in a wrong way.. Sorry for the trouble i will do some more research and ask thanks for your time too

OpenStudy (anonymous):

<.< ask some more lol

OpenStudy (anonymous):

wow you all or just grate

OpenStudy (akshay_budhkar):

is sin(x+y+z) a function?

OpenStudy (anonymous):

hmm... if you say "its a function from R^3 to R" then yes :P

OpenStudy (anonymous):

because you need 3 inputs for it to great one output.

OpenStudy (anonymous):

its a f(x,y,z) = k or whatever 4th variable you want to use lol

OpenStudy (anonymous):

to create* i dont know why i said great lol

OpenStudy (akshay_budhkar):

Oh yes! now i get it! 3 inputs to 1 output makes it a R^3-R function right?

OpenStudy (anonymous):

right.

OpenStudy (akshay_budhkar):

well yes yes i got it!!

OpenStudy (anonymous):

an example of a function from R^3 to R we use all the time is volume. V(l,w,h) = l*w*h three inputs, a one dimensional output.